OpenStudy (anonymous):

bay city and vernonville are 184 mi. apart. a car leaves bay city traveling towards vernonville, and another car leaves vernonville at the same time,traveling bay city. the car leaving bay city avg. 10 mph more than the other, and they meet after 1 hr. and 36 min. what are the avg speeds of the cars

OpenStudy (anonymous):

first establish the variables v = speed of car leaving bay city and v-10 is speed of ther other car baycity car travels x miles and other car travels 184- x miles time taken for each car is 1.6 hours so v = x / 1.6 and v-10 = (184 - x) / 1.6 solve these system of equations and you have your answer

OpenStudy (anonymous):

sorry i am doin it on scratch paper right now just a lil slow

OpenStudy (anonymous):

is there another way of writing out the equation?

OpenStudy (anonymous):

thats ok - i'm getting slow too - age i'm afraid - lol

OpenStudy (anonymous):

x = 1.6v v-10 = (184 - 1.6v) / 1.6 1.6v - 16 = 184 - 1.6v 3.2v = 200 v = 200/3.2 = 62.5 mph - car leaving bay city other car travels at 52.5 mph

OpenStudy (anonymous):

thanks you I appreciate it i got hung up on the equation but you cleared it up thanks!